My question is on linear systems : Angie is twice as old as Amelie. One year ago the sum of their ages was 10. How old is each now? I tried a couple different systems like 2y=x and x+y = 10 but I kept trying to work it out with others to but I can't get it. Please help me! Thank you!
Answer by ankor@dixie-net.com(22740)

Angie is twice as old as Amelie. One year ago the sum of their ages was 10.
How old is each now?
:
Assume you had:
x = Angie
y = Amelie
:
"Angie is twice as old as Amelie."
x = 2y
:
"One year ago the sum of their ages was 10."
(x-1) + (y-1) = 10
x + y - 2 = 10
x + y = 10 + 2
x + y = 12
from the 1st equation, replace x with 2y
2y + y = 12
3y = 12
y = 12/3
y = 4 yrs is Amelie's present age, then obviously
x = 2(4) = 8 yrs is Angies age
;
:
See if this checks out in the statement:
" One year ago the sum of their ages was 10. "
8-1 + 4-1 = 10



How old is each now? 8 and 4